James is five years older than his twin sisters. The sum of the ages of the twins three years ago is equal to his age five years from now. How old will they be 12years from now?
Found 2 solutions by TimothyLamb, ankor@dixie-net.com:
Answer by TimothyLamb(4379)

j = s + 5
j + 5 = 2(s - 3)
j + 5 = 2s - 6
j = 2s - 11
---
j = 2s - 11
s + 5 = 2s - 11
---
s = 16
j = 21
---
How old will they be 12years from now?
Answer:
james = 21 + 12 = 33 years old
each twin sister = 16 + 12 = 28 years old
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by ankor@dixie-net.com(22740)

James is five years older than his twin sisters who are, a & b
j = a + 5
or
a = j-5
and
j = b + 5
or
b = j-5
The sum of the ages of the twins three years ago is equal to his age five years from now.
(a-3) + (b-3) = j + 5
Replace a and b
(j-5) - 3 = (j-5) - 3 = j + 5
2j - 16 = j + 5
2j - j = 5 + 16
j = 21 yrs is James present age
then the twins are 5 yrs younger
a = 16
b = 16
:
How old will they be 12years from now?
Add 12 to the above ages